OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 6y = 36? slope = -2; y-intercept = (0, 36) slope = -2; y-intercept = (0, 6) slope = 6; y-intercept = (0, negative 1 over 3) slope = negative 1 over 3; y-intercept = (0, 6)

OpenStudy (anonymous):

the last one is the correct answer

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

solve for y y=6-(1/2)x using the form y=b+mx whr m= the slope of the line